# LayerTracer: Cognitive-Aligned Layered SVG Synthesis via Diffusion Transformer

### v1

@@ Abstract.
Generating cognitive-aligned layered SVGs remains challenging due to existing methods’ tendencies toward either oversimplified single-layer outputs or optimization-induced shape redundancies. We propose LayerTracer, a diffusion transformer based framework that bridges this gap by learning designers’ layered SVG creation processes from a novel dataset of sequential design operations. Our approach operates in two phases: First, a text-conditioned DiT generates multi-phase rasterized construction blueprints that simulate human design workflows. Second, layer-wise vectorization with path deduplication produces clean, editable SVGs. For image vectorization, we introduce a conditional diffusion mechanism that encodes reference images into latent tokens, guiding hierarchical reconstruction while preserving structural integrity. Extensive experiments demonstrate LayerTracer’s superior performance against optimization-based and neural baselines in both generation quality and editability, effectively aligning AI-generated vectors with professional design cognition. Code is released at https://github.com/showlab/LayerTracer
Diffusion Transformer, Image generation, Vectorization
ccs: Computing methodologies Computer vision
1. Introduction
@@
Scalable Vector Graphics (SVG) serves as the cornerstone of modern digital design, defining visual elements—paths, curves, and geometric shapes—through mathematical equations rather than pixel grids. Unlike raster images, SVG maintains resolution-independent clarity at any scale, making it indispensable for precision-critical applications ranging from UI/UX design to industrial CAD systems. Layered SVGs, which adhere to professional standards, elevate this advantage further: designers can meticulously manipulate individual layers to refine stroke textures, spatial hierarchies, and composite effects. This editability transcends mere convenience—it forms the backbone of dynamic adjustments and collaborative iteration in contemporary design workflows.
@@
Nevertheless, a significant gap persists between current deep learning-based SVG generation techniques and professional requirements. Existing approaches face three systemic challenges: First, the scarcity of large-scale layered SVG datasets forces models to rely on synthetic or oversimplified training data, resulting in outputs devoid of the nuanced hierarchical structures inherent to human designs. Second, methodological fragmentation prevails—optimization-based methods (Jain et al., 2023; Xing et al., 2024b, 2023; Ma et al., 2022; Li et al., 2020; Hirschorn et al., 2024; Wang et al., 2024b) generate vector paths using raster priors but often produce cluttered geometries with redundant anchor points; and large language models (LLMs) (Xing et al., 2024a; Tang et al., 2024; Rodriguez et al., 2023), constrained by token limits, remain limited to basic icons. Most critically, no existing method addresses the designer’s cognitive process—the logical sequencing, spatial reasoning, and element grouping strategies employed during layer construction—resulting in AI-generated SVGs that resemble fragmented collages rather than intentionally editable professional designs.
@@
2. Related Works
2.1. Text2image Diffusion Model
@@
Recent studies have demonstrated that diffusion models are capable of generating high-quality synthetic images, effectively balancing diversity and fidelity. Models based on diffusion models or their variants, such as those paper in (Rombach et al., 2022; Peebles and Xie, 2023), have successfully addressed the challenges associated with text-conditioned image synthesis. Stable Diffusion (Rombach et al., 2022), a model based on the Latent Diffusion Model, incorporates text conditioning within a UNet framework to facilitate text-based image generation (Song et al., 2024b, d, c), establishing itself as a mainstream model in image generation. Fine-tuning pre-trained image generation models can enhance their adaptation to specific application scenarios, as seen in techniques like LoRA (Hu et al., 2021) and DreamBooth (Poole et al., 2022). For theme control in text-to-image generation, several works (Ye et al., 2023; Wang et al., 2024a; Zhang et al., 2024a, b, c, d) focus on custom generation for defined pictorial concepts, with ControlNet (Zhang et al., 2023) additionally offering control over other modalities such as depth information. AnimateDiff (animatediff) introduces a temporal attention module, extending Stable Diffusion into a video generation model. Inspired by ProcessPainter (Song et al., 2024a), which first proposed learning an artist’s painting process through pre-trained temporal models, this paper leverages the in-context capabilities of DiT to generate a layered SVG creation process.
Figure 2. The LayerTracer architecture comprises three key components: (1) Layer-wise Model: Pretrained on our proposed dataset to generate layered pixel sequences from text prompt; (2) Image2Layers Model: Merges LoRA with the Flux base DiT, enabling image-conditioned generation through VAE-encoded latent tokens; (3) Layer-wise Vectorization: Converts raster sequences to SVGs via differential analysis between adjacent layers, followed by Bézier optimization using vtracer to eliminate redundant paths while preserving structural fidelity.
2.2. SVG Generation
@@
Scalable Vector Graphics (SVGs) are widely utilized in design owing to their advantages like geometric manipulability, resolution independence, and compact file structure. SVG generation often involves training neural networks to produce predefined SVG commands and attributes using architectures such as RNNs (Reddy et al., 2021), VAEs (Carlier et al., 2020; Lopes et al., 2019; Tang et al., 2024), and Transformers (Carlier et al., 2020; Wang and Lian, 2021; Wu et al., 2023). Nonetheless, the absence of large-scale vector datasets constrains their generalization capabilities and the creation of complex graphics, with most datasets focusing on specific areas like monochromatic vector icons (Wu et al., 2023) and fonts (Wang and Lian, 2021; Song and Zhang, 2022).
@@
An alternative to directly training an SVG generation network is optimizing it to match a target image during the evaluation phase, employing differentiable rasterizers to bridge vector graphics and raster images (Li et al., 2020). This method optimizes SVG parameters based on pretrained vision-language models. Advances in models like CLIP (Radford et al., 2021) have facilitated effective SVG generation methods such as CLIPDraw (Frans et al., 2021), CLIPasso (Vinker et al., 2022), and CLIPVG (Song et al., 2023). while DreamFusion (Poole et al., 2022) demonstrates the superior generative capabilities of diffusion models. VectorFusion (Jain et al., 2023), DiffSketcher (Xing et al., 2023), and SVGDreamer combine differentiable rasterizers with text-to-image diffusion models to produce vector graphics, achieving notable results in iconography and sketching. However, these methods still face challenges with editability and graphical quality. Recent studies (Thamizharasan et al., 2024; Zhang et al., 2024e) have blended optimization-based methods with neural networks to enhance vector representations by integrating geometric constraints.
@@
2.3. Vectorization
Raster image vectorization or image tracing is a well-studied problem in computer graphics(Bloomenthal and Shoemake, 1991; Carlier et al., 2020; Cook, 1986; Egiazarian et al., 2020). Diffvg(Li et al., 2020) proposes a differentiable rendering method for vectorization, which found shape gradients by differentiating the formula of Reynolds transport theorem with Monta-Carlo edge sampling. Meanwhile, combining differentiable rendering techniques with deep learning models are also studied for image vectorization(Ma et al., 2022; Vinker et al., 2022; Song, 2022). Direct raster-to-vector conversion with neural networks are supported for the relatively simple images(Lopes et al., 2019; Carlier et al., 2020; Reddy et al., 2021). Stroke-based rendering can be used to fit a complex image with a sequence of vector strokes (Singh et al., 2022; Liu et al., 2021; Hu et al., 2023), but the performance is limited by the predefined strokes. Diffvg(Li et al., 2020) can also be leveraged to fit an input image with a set of randomly initialized vector graphical elements. Based on Diffvg, LIVE (Ma et al., 2022) proposes a coarse-to-fine vectorization strategy, with cost tens of minute. CLIPVG(Song et al., 2023) proposes a multi-round vectorization strategy, providing additional graphic elements for the image manipulation task. LIVE (Ma et al., 2022) and O&R (Hirschorn et al., 2024) achieve hierarchical vectorization through optimization-based methods, but their results show a significant gap compared to human-designed works, lacking logical coherence. In contrast to these approaches, our proposed LayerTracer leverages the prior knowledge of the Diffusion Transformer model, reformulating the hierarchical vectorization task as a problem of predicting preceding frames from a reference image.
3. Method
In this section, we begin by exploring the preliminaries on diffusion transformer as detailed in section 3.1. Then introduce the overall architecture of our method in section 3.2, followed by detailed descriptions of the key modules: dataset construction methods in 3.3, Layer-wise image generation in section 3.4, Image Condition Model in section 3.5, and Layer-Wise Vectorization in section 3.6.
3.1. Preliminary
The Diffusion Transformer (DiT) model (Peebles and Xie, 2023), which appears in frameworks such as FLUX.1 (Labs, 2023), Stable Diffusion 3 (Rombach et al., 2022), and PixArt (Chen et al., 2023), employs a transformer-based denoising network to iteratively refine noisy image tokens.
@@ 3.1. Preliminary
In FLUX.1, each DiT block applies layer normalization before Multi-Modal Attention (MMA) (Pan et al., 2020), incorporating Rotary Position Embedding (RoPE) (Su et al., 2024) to capture spatial context. For image tokens $X$, RoPE applies rotation matrices based on a token’s position $(i,j)$ in the 2D grid:
(1) $X_{i,j}\rightarrow X_{i,j}\cdot R(i,j),$
@@
(2) $\text{MMA}([X;C_{T}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$
where $[X;C_{T}]$ denotes concatenation of image and text tokens. This formulation ensures bidirectional attention among the tokens.
3.2. Overall Architecture.
LayerTracer consists of the following components: Serpentine dataset construction, Layer-wise model training, Image2Layers model, and the layer-wise vectorization. Initially, we collected the processes by which designers create layered vector graphics, arranging them in a serpentine layout to form 3x3 and 2x2 grid datasets. Following this, we utilized the LoRA method for pre-training on the proposed dataset, thereby enabling the generation of layered pixel images from textual descriptions. Subsequently, we integrated the LoRA from the previous step with the Flux base model to establish a new foundational model. The Image2Layer model introduces an image-based conditional mechanism that, through additional LoRA fine-tuning, predicts the creation process of reference images. Finally, in the layer-wise vectorization stage, the model sequentially transforms the generated pixel images into high-quality vector graphics, which are analyzed, filtered, and vectorized based on the differences between adjacent layers.
3.3. Serpentine Dataset Construction
Our dataset construction includes 20,000 layered SVGs created by designers, encompassing black outline icons, regular icons, emojis, and illustrative graphics. Each sequence is composed of either 9 or 4 frames, arranged in 3x3 or 2x2 grids, resulting in resolutions of 1056x1056 and 1024x1024 respectively. To capture the process of designers creating layered SVGs, we propose a automated data generation pipeline that deconstructs the layered SVG graphics into sequences based on the grouping logic and element hierarchy within the SVG files. Additionally, the pipeline incorporates a human-in-the-loop process to filter out nonsensical sequences.
In the attention mechanism of DiT, tokens tend to focus on spatially adjacent tokens. This tendency stems from the strong correlations between adjacent image pixels captured during the pre-training of diffusion models. To enhance the model’s learning of grid sequences (Wan et al., 2024), we introduce the serpentine dataset construction method. As shown in Figure 2, we arrange the sequences of 9 and 4 frames in a serpentine layout within the grid, ensuring that temporally adjacent frames are also spatially adjacent (either horizontally or vertically). In our ablation experiments, we confirmed that this design is crucial for the coherence of sequence generation. To facilitate subsequent hierarchical vectorization, for icons with black line strokes, we place the black line layer separately in the first frame during dataset creation. Similarly, during the generation phase, we vectorize the black line layer and overlay it onto the subsequent results.
3.4. Layer-wise Image Generation
@@
(4) $L_{CFM}=E_{t,p_{t}(X|\epsilon),p(\epsilon)}\left[\left\|v_{\Theta}(X,t)-u_{t}(X|\epsilon)\right\|^{2}\right]$
Where $v_{\Theta}(X,t)$ represents the velocity field parameterized by the neural network’s weights,t is timestep, $u_{t}(X|\epsilon)$ is the conditional vector field generated by the model to map the probabilistic path between the noise and true data distributions.
3.5. Image2Layers Model
@@
The primary challenge in training the Image2Layers model lies in the limited availability of high-quality sequential data. While a small dataset may suffice for LoRA training, initializing a controllable plugin (such as ControlNet (Zhang et al., 2023) or IP-Adapter (Ye et al., 2023)) from scratch with limited data is highly challenging. To address this, we design an efficient controllability framework by repurposing a pre-trained DiT model and adapting it to accept image context as a conditioning input. Specifically:
@@
(5) $\text{MMA}([X;C_{I};C_{T}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$
where $[X;C_{I};C_{T}]$ denotes the concatenation of image and text tokens. This formulation enables bidirectional attention.
@@
Figure 3. Given a text prompt, LayerTracer generates cognitive-aligned layered SVGs that mimic human design cognition.
3.6. Layer-Wise Vectorization
@@
Figure 4. Given a raster image of an icon as input, LayerTracer predicts how the icon was created layer by layer, achieving cognitive-aligned layered vectorization.
4. Experiment
4.1. Experiment Setting
Experiment Details. During the pretraining stage, we utilized the Flux 1.0 dev model based on the pretrained DiT architecture. Training resolutions included 1056×1056 (3×3 grids) and 1024×1024 (2×2 grids). The LoRA fine-tuning approach was applied with a LoRA rank of 256, a batch size of 16, and 20,000 fine-tuning steps. For training the Image Condition Model, we merged the Layer-wise Image Generation LoRA with the base model, setting the LoRA merge weight to 1.0. Subsequently, the model was fine-tuned on the same dataset using the LoRA approach for an additional 20,000 steps.
Baseline Methods. The baseline methods in text-to-svg genration methods are SVGDreamer (Xing et al., 2024b), Vecfusion (Jain et al., 2023)and DiffSketcher (Xing et al., 2023). The baseline methods in vectorization include diffvg (Li et al., 2020), LIVE (Ma et al., 2022), and O&R (Hirschorn et al., 2024).
Benchmarks. To address the lack of high-quality layered SVG datasets, this paper introduces a dataset containing over 20,000 layered SVGs and their creation processes, named the LayerSVG Dataset. To ensure fairness in comparative experiments, the Noto-Emoji (Google, 2014) dataset is also included in the benchmark for quantitative evaluation. For the text-to-SVG task and the layer-wise vectorization task, we select 50 prompts and 50 images, respectively, as benchmarks for testing.
4.2. Generation Results
Fig. 3 demonstrates LayerTracer’s capability to generate cognitively-aligned layered SVGs that adhere to text descriptions while maintaining logical layer hierarchies (e.g., background-to-foreground ordering and grouped semantic elements). The outputs preserve essential design properties including layer independence, non-overlapping paths, and topological editability. Fig. 4 further illustrates layer-aware vectorization results, where input raster images are decomposed into clean vector layers with consistent spatial alignment and minimal shape redundancy.
4.3. Comparison and Evaluation

### v3

@@ Abstract
Generating cognitive-aligned layered SVGs remains challenging due to existing methods’ tendencies toward either oversimplified single-layer outputs or optimization-induced shape redundancies. We propose LayerTracer, a DiT based framework that bridges this gap by learning designers’ layered SVG creation processes from a novel dataset of sequential design operations. Our approach operates in two phases: First, a text-conditioned DiT generates multi-phase rasterized construction blueprints that simulate human design workflows. Second, layer-wise vectorization with path deduplication produces clean, editable SVGs. For image vectorization, we introduce a conditional diffusion mechanism that encodes reference images into latent tokens, guiding hierarchical reconstruction while preserving structural integrity. Extensive experiments show that LayerTracer surpasses optimization-based and neural baselines in generation quality and editability. Code is released at https://github.com/showlab/LayerTracer.
{strip}[Uncaptioned image]
Figure 1: LayerTracer creates cognitively aligned layered SVGs from text prompts or by converting images into layered SVGs.
1 Introduction
@@
Scalable Vector Graphics (SVG) is widely used in modern digital design, representing visual elements such as paths, curves, and geometric shapes through mathematical equations rather than pixel grids. Unlike raster images, SVG preserves resolution-independent clarity at any scale, making it suitable for applications requiring high precision, including UI/UX design and industrial CAD systems. Layered SVGs further enhance this flexibility by allowing designers to manipulate individual layers to adjust stroke properties, spatial arrangements, and compositing effects. This structured editability supports dynamic modifications and collaborative workflows in contemporary design practices.
@@
Nevertheless, a significant gap persists between current deep learning-based SVG generation techniques and professional requirements. Existing approaches face three systemic challenges: First, the scarcity of large-scale layered SVG datasets forces models to rely on synthetic or oversimplified training data, resulting in outputs devoid of the nuanced hierarchical structures inherent to human designs. Second, methodological fragmentation prevails—optimization-based methods [18, 58, 57, 24, 20, 14, 51] generate vector paths using raster priors but often produce cluttered geometries with redundant anchor points; and large language models (LLMs) [55, 45, 30, 56], constrained by token limits, remain limited to basic icons . Most critically, no existing method addresses the designer’s cognitive process—the logical sequencing, spatial reasoning, and element grouping strategies employed during layer construction—resulting in AI-generated SVGs that resemble fragmented collages rather than intentionally editable professional designs.
@@
2 Related Works
2.1 Text2image Diffusion Model
@@
Recent studies have demonstrated that diffusion models are capable of generating high-quality synthetic images, effectively balancing diversity and fidelity. Models based on diffusion models or their variants, such as those paper in [31, 26], have successfully addressed the challenges associated with text-conditioned image synthesis. Stable Diffusion [31], a model based on the Latent Diffusion Model, incorporates text conditioning within a UNet framework to facilitate text-based image generation [39, 42, 40], establishing itself as a mainstream model in image generation. Fine-tuning pre-trained image generation models can enhance their adaptation to specific application scenarios, as seen in techniques like LoRA [15] and DreamBooth [27]. For theme control in text-to-image generation, several works [59, 49, 62, 63, 64, 66, 5] focus on custom generation for defined pictorial concepts, with ControlNet [60] additionally offering control over other modalities such as depth information. Some methods [52, 65, 17, 38, 12, 33, 23, 10, 32, 37] explore controllable generation using Diffusion Transformer-based approaches. AnimateDiff [13] introduces a temporal attention module, extending Stable Diffusion into a video generation model. Inspired by ProcessPainter [36], which first proposed learning an artist’s painting process through pre-trained temporal models, this paper leverages the in-context capabilities of DiT to generate a layered SVG creation process.
2.2 SVG Generation
@@
Scalable Vector Graphics (SVGs) are widely utilized in design owing to their advantages like geometric manipulability, resolution independence, and compact file structure. SVG generation often involves training neural networks to produce predefined SVG commands and attributes using architectures such as RNNs [29], VAEs [3, 22, 45], and Transformers [3, 50, 53]. Nonetheless, the absence of large-scale vector datasets constrains their generalization capabilities and the creation of complex graphics, with most datasets focusing on specific areas like monochromatic vector icons [53] and fonts [50, 43].
@@
An alternative to directly training an SVG generation network is optimizing it to match a target image during the evaluation phase, employing differentiable rasterizers to bridge vector graphics and raster images [20]. This method optimizes SVG parameters based on pretrained vision-language models. Advances in models like CLIP [28] have facilitated effective SVG generation methods such as CLIPDraw [9], CLIPasso [47], and CLIPVG [41]. while DreamFusion [27] demonstrates the superior generative capabilities of diffusion models. VecFusion [18], DiffSketcher [57], and SVGDreamer combine differentiable rasterizers with text-to-image diffusion models to produce vector graphics, achieving notable results in iconography and sketching. However, these methods still face challenges with editability and graphical quality. Recent studies [46, 61] have blended optimization-based methods with neural networks to enhance vector representations by integrating geometric constraints.
@@
2.3 Vectorization
Raster image vectorization or image tracing is a well-studied problem in computer graphics[2, 3, 6, 8]. Diffvg[20] proposes a differentiable rendering method for vectorization, which found shape gradients by differentiating the formula of Reynolds transport theorem with Monta-Carlo edge sampling. Meanwhile, combining differentiable rendering techniques with deep learning models are also studied for image vectorization[24, 47, 35]. Direct raster-to-vector conversion with neural networks are supported for the relatively simple images[22, 3, 29]. Stroke-based rendering can be used to fit a complex image with a sequence of vector strokes [34, 21, 16], but the performance is limited by the predefined strokes. Diffvg[20] can also be leveraged to fit an input image with a set of randomly initialized vector graphical elements. Based on Diffvg, LIVE [24] proposes a coarse-to-fine vectorization strategy, with cost tens of minute. CLIPVG[41] proposes a multi-round vectorization strategy, providing additional graphic elements for the image manipulation task. LIVE [24] and O&R [14] achieve hierarchical vectorization through optimization-based methods, but their results show a significant gap compared to human-designed works, lacking logical coherence. In contrast to these approaches, our proposed LayerTracer leverages the prior knowledge of the Diffusion Transformer model, reformulating the hierarchical vectorization task as a problem of predicting preceding frames from a reference image.
Figure 2: The LayerTracer architecture comprises three key components: (1) Layer-wise Model: Pretrained on our proposed dataset to generate layered pixel sequences from text prompt; (2) Image2Layers Model: Merges LoRA with the Flux base DiT, enabling image-conditioned generation through VAE-encoded latent tokens; (3) Layer-wise Vectorization: Converts raster sequences to SVGs via differential analysis between adjacent layers, followed by Bézier optimization using vtracer to eliminate redundant paths while preserving structural fidelity.
3 Method
In this section, we begin by exploring the preliminaries on diffusion transformer as detailed in section 3.1. Then introduce the overall architecture of our method in section 3.2, followed by detailed descriptions of the key modules: dataset construction methods in 3.3, Layer-wise image generation in section 3.4, Image2Layers in section 3.5, and Layer-Wise Vectorization in section 3.6.
3.1 Preliminary
The Diffusion Transformer (DiT) model [26], which appears in frameworks such as FLUX.1 [19], Stable Diffusion 3 [31], and PixArt [4], employs a transformer-based denoising network to iteratively refine noisy image tokens.
@@ 3.1 Preliminary
In FLUX.1, each DiT block applies layer normalization before Multi-Modal Attention (MMA) [25], incorporating Rotary Position Embedding (RoPE) [44] to capture spatial context. For image tokens $X$, RoPE applies rotation matrices based on a token’s position $(i,j)$ in the 2D grid:
$X_{i,j}\rightarrow X_{i,j}\cdot R(i,j),$ (1)
@@
$\text{MMA}([X;C_{T}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (2)
where $X;C_{T}]$ denotes concatenation of image and text tokens. This formulation ensures bidirectional attention among the tokens.
3.2 Overall Architecture.
LayerTracer consists of the following components: Serpentine dataset construction, Layer-wise model training, Image2Layers model, and the layer-wise vectorization. Initially, we collected the processes designers use to create layered vector graphics and organized them into grid-based datasets. Following this, we utilized the LoRA method for pre-training on the proposed dataset, thereby enabling the generation of layered pixel images from textual descriptions. Subsequently, we integrated the LoRA from the previous step with the Flux base model to establish a new foundational model. The Image2Layers model introduces an image-based conditional mechanism that, through additional LoRA fine-tuning, predicts the creation process of reference images. Finally, in the layer-wise vectorization stage, the model sequentially transforms the generated pixel images into high-quality vector graphics, which are analyzed, filtered, and vectorized based on the differences between adjacent layers.
3.3 Serpentine Dataset Construction
Our dataset construction includes 20,000 layered SVGs created by designers, encompassing black outline icons, regular icons, emojis, and illustrative graphics. As shown in Fig. 2, each sequence is composed of either 9 or 4 frames, arranged in 3x3 or 2x2 grids, resulting in resolutions of 1056x1056 and 1024x1024 respectively. To capture the process of designers creating layered SVGs, we propose a automated data generation pipeline that deconstructs the layered SVG graphics into sequences based on the grouping logic and element hierarchy within the SVG files. Additionally, the pipeline incorporates a human-in-the-loop process to filter out nonsensical sequences.
In the attention mechanism of DiT, tokens tend to focus on spatially adjacent tokens. This tendency stems from the strong correlations between adjacent image pixels captured during the pre-training of diffusion models. To enhance the model’s learning of grid sequences [48], we introduce the serpentine dataset construction. As shown in Figure 2, we arrange the sequences of 9 and 4 frames in a serpentine layout within the grid, ensuring that temporally adjacent frames are also spatially adjacent (either horizontally or vertically). In our ablation experiments, we confirmed that this design is crucial for the coherence of sequence generation. To aid hierarchical vectorization, icons with black strokes isolate the line layer in the first frame during dataset creation. During the generation phase, we vectorize the black line layer and overlay it onto the subsequent results.
3.4 Layer-wise Image Generation
@@
$L_{\mathrm{CFM}}=\mathbb{E}_{t,\;p_{t}(X|\epsilon),\;p(\epsilon)}\!\bigl{[}\bigl{\|}\,v_{\Theta}(X,t)-u_{t}(X|\epsilon)\bigr{\|}^{2}\bigr{]}$ (4)
Where $v_{\Theta}(X,t)$ represents the velocity field parameterized by neural network weights, $t$ denotes the timestep, and $u_{t}(X|\epsilon)$ is the conditional vector field mapping the path between noise and true data distributions.
3.5 Image2Layers Model
@@
The primary challenge in training the Image2Layers model lies in the limited availability of high-quality sequential data. While a small dataset may suffice for LoRA training, initializing a controllable plugin (such as ControlNet [60] or IP-Adapter [59]) from scratch with limited data is highly challenging. To address this, we design an efficient controllability framework by repurposing a pre-trained DiT model and adapting it to accept image context as a conditioning input. Specifically:
@@
$\text{MMA}([X;C_{I};C_{T}])=\text{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (5)
where $X;C_{I};C_{T}]$ denotes the concatenation of image and text tokens. This formulation enables bidirectional attention.
@@
Figure 3: Given a text prompt, LayerTracer generates cognitive-aligned layered SVGs that mimic human design cognition.
3.6 Layer-Wise Vectorization
@@
Figure 4: Given a raster image of an icon as input, LayerTracer predicts how the icon was created layer by layer, achieving cognitive-aligned layered vectorization.
4 Experiment
4.1 Experiment Setting
Experiment Details. During the pretraining stage, we utilized the Flux 1.0 dev model based on the pretrained DiT architecture. Training resolutions included 1056×1056 (3×3 grids) and 1024×1024 (2×2 grids). The LoRA fine-tuning approach was applied with a LoRA rank of 256, a batch size of 16, a learning rate of 0.001, and 20,000 fine-tuning steps. For the three styles of SVGs, we trained separate LoRA models for black outline icons, illustrations, and emojis, using 3K, 2K, and 15K training samples, respectively. During the Image2Layers Model training phase, we merged the LayerTracer LoRA with the base model, using a LoRA merge weight of 1.0. Then, we fine-tuned it for 20,000 steps using LoRA-based training on the same dataset as the Layer-wise Model training. All training was conducted on a single A100 GPU with 80GB of memory.
Baseline Methods. The baseline methods in text-to-svg genration methods are SVGDreamer [58], Vecfusion [18]and DiffSketcher [57]. The baseline methods in vectorization include diffvg [20], LIVE [24], and O&R [14]. All baseline methods are evaluated with their default settings to ensure fairness.
Benchmarks. To address the lack of high-quality layered SVG datasets, this paper introduces a dataset containing over 20,000 layered SVGs and their creation processes, named the LayerSVG Dataset. To ensure fairness in comparative experiments, the Noto-Emoji [11] dataset is also included in the benchmark for quantitative evaluation. For the text-to-SVG task and the layer-wise vectorization task, we select 50 prompts and 50 images, respectively, as benchmarks for testing.
4.2 Generation Results
Figure 5: Layer-wise SVG generation with color gradients.
Fig. 3 demonstrates LayerTracer’s capability to generate cognitively-aligned layered SVGs that adhere to text descriptions while maintaining logical layer hierarchies (e.g., background-to-foreground ordering and grouped semantic elements). The outputs preserve essential design properties including layer independence, non-overlapping paths, and topological editability. Fig. 4 further illustrates layer-aware vectorization results, where input raster images are decomposed into clean vector layers with consistent spatial alignment and minimal shape redundancy. Fig. 5 shows that training with gradient-colored samples and advanced vectorization methods [7] enables LayerTracer to generate layered vector graphics with color gradients. More experimental results can be found in the supplementary material.
4.3 Comparison and Evaluation